OpenStudy (abbycross167):

Can someone please help me with this math question? Please? I'll fan and give a medal!!

OpenStudy (abbycross167):

Find the inverse of x^2/2. Is the inverse a function? Explain

jimthompson5910 (jim_thompson5910):

\[\Large f(x) = \frac{x^2}{2}\] \[\Large y = \frac{x^2}{2}\] \[\Large x = \frac{y^2}{2}\] now solve for y

OpenStudy (abbycross167):

Y=\/2x @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you mean \[\Large y = \sqrt{2x}\] right?

OpenStudy (abbycross167):

Yes sir

jimthompson5910 (jim_thompson5910):

you forgot about the fact that if x^2 = 25 then x = 5 or x = -5

jimthompson5910 (jim_thompson5910):

so there should be a plus/minus \[\Large y = \pm \sqrt{2x}\]

OpenStudy (abbycross167):

So I could write it like this? Y=\/2x or Y=-\/2x

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

you can write sqrt(x) to signify \(\Large \sqrt{x}\)

jimthompson5910 (jim_thompson5910):

graph both of the equations together. Is it possible to pass a vertical line through more than one point on the graph?

jimthompson5910 (jim_thompson5910):

think of the two curves as one big curve

OpenStudy (abbycross167):

yes a vertical line can

jimthompson5910 (jim_thompson5910):

so it fails the vertical line test, meaning that the inverse is NOT a function

OpenStudy (abbycross167):

Okay thank you very much for helping me!!

jimthompson5910 (jim_thompson5910):

np